OpenStudy (anonymous):

For questions 22–23, solve the equation by completing the square. Round to the nearest hundredth if necessary. 20. x² + 10x = 18 21. x² + 9x – 14 = 0

OpenStudy (anonymous):

Ok I can do this

OpenStudy (snuggielad):

answering history and bio questions today...passing it back @Luis_Rivera

OpenStudy (anonymous):

20. x² + 10x = 18 (1 point) • –11.56, 1.56 ** • 11.56, 1.56 • –11.56, –1.56 • 11.56, –1.56 21. x² + 9x – 14 = 0 (1 point) • 10.35, 1.35 • 10.35, –1.35 • –10.35, –1.35 • –10.35, 1.35**

OpenStudy (anonymous):

Thank you Just 1 more

OpenStudy (anonymous):

Ok np

OpenStudy (anonymous):

Hit me